Shift Summary for Eve Shift, 12/06/99

Patricia McBride (mcbride@fnpsph)
Mon, 06 Dec 1999 23:26:01 CST

Shift Summary for Eve shift, 12/06/99

On Shift: P. McBride

Others present: John Shields, Peter Shanahan, Sydney, Jason Hamm, Ashkan,
Valeri, Sasha L.

========================================================================

Runs ended this shift:

15143 -- LASER_SCAN -- ET cable swapped
15145 -- LASER_SCAN -- dynode cable swapped
15146 -- COSMICS to disk -- all cables are back in normal locations
run continues after shift crew goes home

========================================================================

Overview of shift:

Continuation of TeV downtime for magnet repairs.

J. Shields worked on trigger 1 skim.
Jason did a laser scan after swapping an pair of ET cable (Run 15143);
ET cables swapped back after scan is completed.
TRD_ALAM -- Sasha L paged... he is investigating.
Ashkan updates SDAQ to run with CERNLIB 2000.
There are problems with the TRD monitoring with new libraries.
Valeri reinstalls the old executables for TRD monitoring.
TRD alarms go away.
Cosmic run started.
P. McBride did the gas check list.
Jason and Peter swapped a dynode cable and did a laser scan (Run 15145).
Julie requested a cosmic run with data written to disk;
cosmic run started (Run 15146).

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shifts

Keep an eye on CsI channel 2446

-- 
Patricia McBride          mcbride@fnal.gov          PHONE: (630)840-8071
Fermi National Accelerator Laboratory               FAX:   (630)840-2783